# KungfuBot: Physics-Based Humanoid Whole-Body Control for Learning Highly-Dynamic Skills

### v1

@@ Appendix A Derivation of Optimal Tracking Sigma
$\max_{\sigma\in\mathbb{R_{+}}}\quad$ $J^{\mathrm{ex}}(\bm{x}^{*})$ (9a)
@@ C.2 Reward Design
Joint position limits $\mathbb{I}(\bm{q}\notin[\bm{q}_{\mathrm{soft\text{-}min}},\ \bm{q}_{\mathrm{soft\text{-}max}}])$10.0
Joint velocity limits $\mathbb{I}(\dot{\bm{q}}\notin[\dot{\bm{q}}_{\rm{soft\text{-}min}},\ \dot{\bm{q}}_{\rm{soft\text{-}max}}])$5.0
Joint torque limits $\mathbb{I}(\bm{\tau}\notin[\bm{\tau}_{\rm{soft\text{-}min}},\ \bm{\tau}_{\rm{soft\text{-}max}}])$5.0
Slippage $\|\bm{v}_{\rm xy}^{\rm feet}\|_{2}^{2}\cdot\mathbb{I}[\|F^{\rm feet}\|_{2}\geq 1]$1.0
Feet contact forces $\min(\|F^{\rm feet}-400\|_{\rm 2}^{\rm 2},0)$0.01
Feet air time[30] $\mathbb{I}[T_{\rm{air}}>0.3]$1.0
Stumble $\mathbb{I}[\left\|\bm{F}_{{xy}}^{\rm{feet}}\right\|>5\cdot F_{z}^{\rm{feet}}]$2.0
Torque $\|\bm{\tau}\|_{2}^{2}$1e-6
Action rate $\|\bm{a}_{t}-\bm{a}_{t-1}\|_{2}^{2}$0.02
Collision $\mathbb{I}_{\rm{collision}}$30
Termination $\mathbb{I}_{\rm{termination}}$200
@@ C.3 Domain Randomization
[…]
ERFI[58](N·m/kg) 0.05×0.05$\times$0.05 × torque limit
[…]
@@
Random push interval(s) [5, 10]510[5,\ 10][ 5 , 10 ]

### v2

@@ Appendix A Derivation of Optimal Tracking Sigma
$\max_{\sigma\in\mathbb{R_{+}}}\qquad$ $J^{\mathrm{ex}}(\bm{x}^{*})$ (9a)
@@ C.2 Reward Design
Joint position limits $\mathbb{I}(\bm{q}\notin[\bm{q}_{\mathrm{soft\text{-}min}},\ \bm{q}_{\mathrm{soft\text{-}max}}])$ -10.0
Joint velocity limits $\mathbb{I}(\dot{\bm{q}}\notin[\dot{\bm{q}}_{\rm{soft\text{-}min}},\ \dot{\bm{q}}_{\rm{soft\text{-}max}}])$ -5.0
Joint torque limits $\mathbb{I}(\bm{\tau}\notin[\bm{\tau}_{\rm{soft\text{-}min}},\ \bm{\tau}_{\rm{soft\text{-}max}}])$ -5.0
Slippage $\|\bm{v}_{\rm xy}^{\rm feet}\|_{2}^{2}\cdot\mathbb{I}[\|F^{\rm feet}\|_{2}\geq 1]$ -1.0
Feet contact forces $\min(\|F^{\rm feet}-400\|_{\rm 2}^{\rm 2},0)$ -0.01
Feet air time[30] $\mathbb{I}[T_{\rm{air}}>0.3]$ -1.0
Stumble $\mathbb{I}[\left\|\bm{F}_{{xy}}^{\rm{feet}}\right\|>5\cdot F_{z}^{\rm{feet}}]$ -2.0
Torque $\|\bm{\tau}\|_{2}^{2}$ -1e-6
Action rate $\|\bm{a}_{t}-\bm{a}_{t-1}\|_{2}^{2}$ -0.02
Collision $\mathbb{I}_{\rm{collision}}$ -30
Termination $\mathbb{I}_{\rm{termination}}$ -200
@@ C.3 Domain Randomization
Friction $\mathcal{U}(0.2,\ 1.2)$
PD gain $\mathcal{U}(0.9,\ 1.1)$
Link mass(kg) $\mathcal{U}(0.9,\ 1.1)\times$ default
Ankle inertia(kg·m2) $\mathcal{U}(0.9,\ 1.1)\times$ default
Base CoM offset(m) $\mathcal{U}(-0.05,\ 0.05)$
ERFI[58](N·m/kg) $0.05\times$ torque limit
Control delay(ms) $\mathcal{U}(0,\ 40)$
@@
Random push interval(s) [5, 10][5,\ 10]